If the demand curve for bread is given as Q = 300 - 16p, (5 marks)
a. What is the point elasticity of demand when price is 1.50?
b. What will be your pricing decision if you want to increase total revenue? To increase
or decrease price?
1
Expert's answer
2021-08-23T13:11:19-0400

a) Point elasticity of demand is given by :-

"-\frac{dQ}{dP}\times \frac{P}{Q}"

Differentiating Q with respect to p we get

"\frac{dQ}{dp} = -16\\

Given \\p = 1.50"

Substituting this value of p in the demand equation we get

"Q=300-16(1.5)=300-24\\
Q=276"

 

Therefore, elasticity

"= - (-16 × \frac{1.50}{276}) = 0.09 (approx.)"


b) Here value of elasticity of demand if less than 1, thus implying that demand is inelastic. This means that even if price rises by a large value, change in quantity demanded will be very small. Hence, in order to increase total revenue appropriate decision would be to increase price. In such a case, quantity demanded won't fall much and as such, the firm will be able to enjoy a higher revenue.